# Channel-Level Calibration Methods of Silicon Photomultiplier for JUNO-TAO Central Detector

### v1

@@ 6.3 Correct Dark Count Rate Base on the External Optical Crosstalk Rate Calibration Result
7 Conclusion & Limitation
7.1 Conclusion
This study presents a systematic investigation of the channel-level and tile-level calibration of key parameters for the silicon photomultiplier (SiPM) and its associated electronics in the TAO central detector. The data sets used in this study are summarized in Table 2.
At the channel level, the calibration approaches for individual parameters are as follows:
  1. 1.
    Dark Count Rate (DCR): Calibrated by counting the number of hits recorded before the trigger time.
  2. 2.
    Time Offset: Calibrated using external trigger LED events with the source positioned at the center of the LS. The time offset for each channel is determined by analyzing the difference in the most probable hit time value for each channel.
  3. 3.
    Relative Photon Detection Efficiency (PDE): Calibrated using $^{68}{\rm Ge}$ events with the source positioned at the center of the liquid scintillator (LS), where the relative PDE between channels is determined by the relative number of zero-hit events across these channels.
  4. 4.
    Gain: The charge spectrum is analyzed using a multi-Gaussian function fit. The gain value is determined by the separation between adjacent Gaussian peaks in the multi-Gaussian function.
  5. 5.
    Internal Optical Crosstalk (IOCT): An initial approach calibrated the IOCT by analyzing the proportion of multi-photoelectrons (multi-PEs) components in DN events. However, this method is susceptible to inaccuracies due to intrinsic multi-PEs DN. To overcome this limitation, an improved method was developed that employs a generalized Poisson distribution fit to the DN charge spectrum, resulting in a more reliable IOCT calibration.
At the tile level, two critical parameters are calibrated:
  1. 1.
    External Optical Crosstalk (EOCT) Rate: The operating state of SiPMs is controlled by using the high-voltage system. The EOCT rate for a target SiPM is determined by measuring the difference in response when all other SiPMs are turned on and turned off.
  2. 2.
    External Optical Crosstalk Emission Angle Distribution: Calibration is performed by selectively turning on/off SiPMs within a specific angular range relative to a target SiPM. The bias is less than 4% within the main angular range of EOCT emission photons.
    The bias and standard deviation of all calibration parameter results are summarized in Table 3.
Table 3: The Bias and Standard Deviation of Each Calibrated Parameter
Calibrated Parameter Bias Standard Deviation
DCR (Uncorrected) 23.6% 1.66%
DCR (Corrected) -0.40% 1.32%
Gain 0.084% 0.094%
Time Offset 0.027 ns 0.031ns
Relative PDE <3% 111between the channel at pole and equator. 0.17%
IOCT(Multi-PEs) 5.70% 2.73%
IOCT(GP Fitting) 1.40% 3.02%
EOCT Rate 0.009% 0.096%
7.2 Limitation
This paper does not present a calibration method for the AP. As shown in Figure 19, AP contributes a fine structure to the tail of each PE peak in the charge spectrum. However, electronic noise and the charge reconstruction algorithm implemented on the FPGA can produce similar fine structures. Since the noise level and charge reconstruction algorithm differ between simulated and real data, and considering that the AP rate measured in the TAO SiPM mass test is very low [Ref32], this paper does not provide a calibration method for AP. A dedicated calibration method will be proposed in future work based on real data.
Figure 19: Charge Spectrum Distributions with 1% AP and without AP. The red line represents the case with AP, while the blue line represents the case without AP. It can be observed that AP contributes a fine structure to the tail of each PE peak in the charge spectrum.
Similarly, due to the electronic environment, the charge and time reconstruction algorithm in reality differ from those used in the simulation. The calibration method presented in this paper will be optimized based on real data in future work, and the corresponding calibration results will be provided subsequently.
This paper does not provide a method for correcting the optical reflection effect on the SiPM surface. A solution will be updated in future work based on Geant4 detector optical simulation.
@@ Acknowledgments
We are gratefully for the JUNO-TAO and the technical staff of the participating institutions. This work was supported by National Natural Science Foundation of China under Grant Number 12275281.

### v2

@@ 7.1 Impact of Temperature Fluctuation on the TAO SiPM Parameters
[…]
Substituting Equation 7.4 into the expression for $\sigma_{T,\text{DCR}}$ in Equation 7.3, the following expression can be derived:
[…]
Parameter Relative Uncertainty
DCR 0.75%
Gain 0.17%
PDE 0.058%
IOCT 0.17%
EOCT Rate 0.17%
7.2 Impact of SiPM Parameter Uncertainties on the Reconstructed Vertex and Energy
In the TAO simulation, we individually add the uncertainties of each SiPM parameter, which originate from temperature fluctuation and the calibration method (the standard deviation derive from calibration). We simulate 1 MeV kinetic energy electrons at the center of the liquid scintillator in the TAO CD, reconstruct the vertex and energy of each electron via the charge center method [39], and derive the corresponding reconstructed vertex uncertainty and energy resolution for both scenarios with and without the inclusion of SiPM parameter uncertainties. The results are summarized in Table 3.
Table 3: Impact of SiPM Parameter Uncertainties on Reconstructed Vertex Uncertainty and Energy Resolution
SiPM Parameter Uncertainty Source Vertex Uncertainty Energy Resolution
Absolute ($\mathrm{m}\mathrm{m}$) Relative ($\%$) Absolute ($\%$) Relative ($\%$)
Baseline No Uncertainty $13.080$ $0.000$ $1.8081$ $0.0000$
DCR Temperature $13.082$ $0.015$ $1.8084$ $0.0166$
PDE $13.082$ $0.015$ $1.8084$ $0.0166$
Gain $13.082$ $0.015$ $1.8104$ $0.1272$
IOCT $13.081$ $0.008$ $1.8083$ $0.0111$
EOCT $13.082$ $0.015$ $1.8086$ $0.0277$
All Parameters $13.089$ $0.0688$ $1.8117$ $0.1991$
DCR Calibration $13.083$ $0.023$ $1.8085$ $0.0221$
PDE $13.099$ $0.145$ $1.8107$ $0.1438$
Gain $13.081$ $0.008$ $1.8089$ $0.0442$
IOCT $13.097$ $0.130$ $1.8134$ $0.2931$
EOCT $13.082$ $0.015$ $1.8084$ $0.0166$
All Parameters $13.121$ $0.313$ $1.8270$ $1.0453$
All Parameters All Uncertainties $13.131$ $0.390$ $1.8300$ $1.2112$
According to the results in Table 3, it can be seen that the SiPM parameter uncertainties induced by temperature fluctuation and those introduced by the calibration method have a very limited impact on the vertex and energy reconstruction precision obtained via the charge center method [39]. The additional vertex smearing and extra degradation of energy resolution introduced by these uncertainties are very small, which directly confirms that the calibration precision fully meets the performance requirements of the TAO detector.
8 Conclusion
This work presents a systematic and comprehensive calibration strategy for the SiPMs in the TAO CD, covering the core performance parameters including DCR, time offset, relative PDE, gain, and IOCT rate. Furthermore, we propose a novel method based on the SiPM on-off switching scheme, which is specifically designed for the calibration of the EOCT rate and its emission angular distribution in detector with high dark noise levels. The feasibility of all proposed calibration methods is fully verified based on the TAO offline software simulation framework, with the bias, standard deviation, and underlying physical sources of bias for each method systematically quantified.
For the key SiPM parameters, the core calibration results are summarized as follows: the DCR calibration achieves a bias of -0.4% and a standard deviation of 1.32% after correction with the EOCT rate calibration results, which effectively suppresses the 23.6% large bias caused by EOCT; the time offset calibration yields both bias and standard deviation better than 0.2 ns, where the uncertainty is dominated by the approximately 1 cm positional deviation of the deployed source inside the detector; for the relative PDE calibration, after correcting for the contribution of dark noise, the method achieves a maximum bias of approximately 3% and a standard deviation of 0.17%, which originates from the non-uniformity of the light field induced by optical reflection on the SiPM surface; the IOCT rate calibration using the Generalized Poisson distribution fitting method achieves a bias of 1.40% and a standard deviation of 3.02%, which resolves the 5.70% bias from dark noise-induced multiple PEs hits, with the residual bias dominated by the contribution from AP.
For the proposed novel EOCT calibration method, the SiPM on-off switching scheme enables EOCT rate calibration with both bias and standard deviation better than 0.1%. Meanwhile, we realize the calibration of the EOCT emission angle distribution by controlling the SiPMs on-off in each angular interval, achieving a bias better than 4% and a standard deviation better than 1% at a 5-degree angular interval.
In addition, this work investigates the SiPM parameter uncertainties induced by temperature fluctuation in the TAO CD. We incorporate the uncertainties from both temperature fluctuation and the calibration method into the TAO simulation, and evaluate their impact on event reconstruction via the charge center method. The results demonstrate that the corresponding impact is very limited, which confirms that the proposed calibration scheme fully meets the stringent performance requirements of the TAO detector.
This work provides solid theoretical guidance for the calibration of individual SiPM performance parameters in the future TAO real detector. Accurate calibration of SiPM parameters will facilitate an in-depth understanding of the TAO detector response and the origins of its energy resolution, and further support the optimization of particle event reconstruction algorithms and the overall performance of the detector.
@@ Acknowledgments
We are gratefully for the JUNO-TAO and the technical staff of the participating institutions. This work is supported by the National Key Research and Development Project of China, Grant No. 2022YFA1602002, National Natural Science Foundation of China under Grant Number 12275281.